### v1

# Worth Their Weight: Randomized and Regularized Block Kaczmarz Algorithms without Preprocessing

Abstract
Due to the ever growing amounts of data leveraged for machine learning and scientific computing, it is increasingly important to develop algorithms that sample only a small portion of the data at a time. In the case of linear least-squares, the randomized block Kaczmarz method (RBK) is an appealing example of such an algorithm, but its convergence is only understood under sampling distributions that require potentially prohibitively expensive preprocessing steps. To address this limitation, we analyze RBK when the data is sampled uniformly, showing that its iterates converge in a Monte Carlo sense to a weighted least-squares solution. Unfortunately, for general problems the condition number of the weight matrix and the variance of the iterates can become arbitrarily large. We resolve these issues by incorporating regularization into the RBK iterations. Numerical experiments, including examples arising from natural gradient optimization, suggest that the regularized algorithm, ReBlocK, outperforms minibatch stochastic gradient descent for realistic problems that exhibit fast singular value decay.
Machine Learning, SciML, Natural Gradient Descent, Linear Regression, Kaczmarz
1 Introduction
Consider the linear least-squares problem
$\begin{split}&\min_{x\in\mathbb{R}^{n}}\left\lVert Ax-b\right\rVert^{2},\\ A&\in\mathbb{R}^{m\times n},\;b\in\mathbb{R}^{m}.\end{split}$ (1)
The minimal-norm ordinary least-squares solution is
$x^{*}=A^{+}b,$ (2)
where $A^{+}$ is the Moore-Penrose pseudoinverse of $A$.
In this work, we are interested in algorithms that solve (1) by sampling just a small number of rows at a time. Such algorithms can be useful for extremely large problems for which direct methods and Krylov subspace methods are prohibitively expensive (Censor, 1981). Of particular interest are problems in which the dimension $n$ is so large that it is only possible to access $k\ll n$ rows at a time, but not so large that it is necessary to take $k=1$. As a canonical example, we might have $m=10^{9}$, $n=10^{6}$, and $k=10^{3}$.
As further motivation for algorithms of this type, there are some applications in which sampling rows is the only efficient way to access the data. For example, the rows may be computed on the fly, especially when solving the “semi-infinite” version of (1) in which the rows are indexed by continuous variables rather than discrete integers (Shustin & Avron, 2022). An example that captures both of these features is the problem of calculating natural gradient directions for continuous function learning problems; see Section 5.
One well-known approach for solving (1) using only a few rows at a time is minibatch stochastic gradient descent (mSGD). The mSGD algorithm works by sampling $k$ rows uniformly at random and using them to calculate an unbiased estimator of the gradient of the least-squares loss function. This is equivalent to averaging the $k$ independent gradient estimators furnished by each sampled row. For a thorough discussion of mSGD, see (Jain et al., 2018b).
The technique of averaging used in mSGD, while simple and efficient, is a relatively crude way of processing a block of $k$ rows. A more sophisticated approach is provided by the randomized block Kaczmarz method, which goes beyond averaging by making use of the pseudoinverse of the sampled block (Needell & Tropp, 2014). As we shall see, the use of the pseudoinverse opens up the possibility of faster convergence, but it also creates a number of complications regarding the implementation and analysis of the algorithm.
1.1 Notation
We denote by $r=b-Ax^{*}$ the residual vector of (1), by $a_{i}^{\top}\in\mathbb{R}^{n}$ the row of $A$ with index $i$, and by $b_{i}\in\mathbb{R}$ the corresponding entry of $b$. Additionally, for an index set $S\subseteq\left\{1,\ldots,m\right\}$ with $|S|=k$, let $A_{S}\in\mathbb{R}^{k\times n}$ represent the block of rows of $A$ whose indices are in $S$, and let $b_{S}\in\mathbb{R}^{k}$ represent the corresponding entries of $b$. The same subscript notations will also be applied as needed to any matrices and vectors other than $A$ and $b$. Additionally, denote by $\mathbf{U}(m,k)$ the uniform distribution over all size-$k$ subsets of $\left\{1,\ldots,m\right\}$.
For symmetric matrices $X,Y$, denote by $X\succ Y$ that the difference $X-Y$ is positive definite, and define ,$\succeq$,$\prec$,$\preceq$ , ≺ , ⪯ correspondingly. For any vector $x$ and any positive definite matrix $Y$ of the same size, let $\left\lVert x\right\rVert_{Y}=\sqrt{x^{\top}Yx}$. For any matrix $X$, denote by $\|X\|_{F}$ its Frobenius norm and by $\sigma_{\rm min}^{+}(X)$ its minimum nonzero singular value.
1.2 Randomized Kaczmarz
The modern version of the randomized Kaczmarz method (RK) was proposed by Thomas Strohmer and Roman Vershynin (2009). In RK, an initial guess $x_{0}$ is updated iteratively using a two-step procedure:
  1. 1.
    Sample a row index $i_{t}\in\left\{1,\ldots,m\right\}$ with probability proportional to $\left\lVert a_{i_{t}}\right\rVert^{2}$.
  2. 2.
    Update
    $x_{t+1}=x_{t}+a_{i_{t}}\frac{b_{i_{t}}-a_{i_{t}}^{\top}x_{t}}{\left\lVert a_{i_{t}}\right\rVert^{2}}.$ (3)
To reduce the sampling cost, it is also possible to run RK with uniform sampling, which is equivalent to running RK on a diagonally reweighted problem (Needell et al., 2014).
The iteration (3) has the interpretation of projecting $x_{t}$ onto the hyperplane of solutions to the individual equation $a_{i_{t}}^{\top}x=b_{i_{t}}$. For consistent systems, $Ax^{*}=b$, the RK iterates $x_{t}$ converge linearly to $x^{*}$ with a rate that depends on the conditioning of $A$. For inconsistent systems, $Ax^{*}\neq b$, RK converges only to within a finite horizon of the ordinary least-squares solution $x^{*}$ (Needell, 2010). In particular, the expected squared error $\mathbb{E}\,\left\lVert x_{t}-x^{*}\right\rVert^{2}$ converges to a finite, nonzero value that depends on the conditioning of $A$ and the norm of the residual vector $r=b-Ax^{*}$. See Theorem 7 of (Zouzias & Freris, 2013) for the strongest known convergence bound of this type.
1.3 Tail Averaging
Tail averaging is a common technique for boosting the accuracy of stochastic algorithms (Rakhlin et al., 2011; Jain et al., 2018b; Epperly et al., 2024). Given a series of stochastic iterates $x_{0},\ldots,x_{T}$ and a burn-in time $T_{b}$, the tail-averaged estimator is given by
$\overline{x}_{T}=\frac{1}{T-T_{b}}\sum_{t=T_{b}+1}^{T}x_{t}.$ (4)
The recent work (Epperly et al., 2024) shows that applying tail averaging to the RK iterates yields exact convergence (with no finite horizon) to the ordinary least-squares solution $x^{*}$, even for inconsistent systems. Building on these results, we will make use of tail averaging to obtain exact convergence to a weighted least-squares solution in the block case.
1.4 Randomized Block Kaczmarz
Randomized block Kaczmarz (RBK) is an extension of RK which uses blocks of rows to accelerate the convergence and make better use of parallel and distributed computing resources (Elfving, 1980; Needell & Tropp, 2014). Like RK, each RBK iteration proceeds in two steps:
  1. 1.
    Sample a subset $S_{t}\subset\left\{1,\ldots,m\right\}$ of the row indices from some chosen sampling distribution $\rho$.
  2. 2.
    Update
    $x_{t+1}=x_{t}+A_{S_{t}}^{+}(b_{S_{t}}-A_{S_{t}}x).$ (5)
Here $A_{S_{t}}^{+}$ is the Moore-Penrose pseudoinverse of $A_{S_{t}}$. The iteration (5) has the interpretation of projecting $x_{t}$ onto the hyperplane of solutions to the block of equations $A_{S_{t}}x=b_{S_{t}}$. The RBK method can also be viewed as a “sketch-and-project” algorithm; see (Gower & Richtárik, 2015).
There have been many proposals for how to choose the blocks in the RBK method. One idea is to use a preprocessing step to partition the matrix $A$ into well-conditioned blocks, then sample this fixed set of blocks uniformly (Needell & Tropp, 2014). It has also been suggested to preprocess the matrix with an incoherence transform, which can make it easier to generate a well-conditioned partition (Needell & Tropp, 2014) or, relatedly, enable RBK with uniform sampling to converge rapidly for the transformed problem (Dereziński & Yang, 2024). The work of (Dereziński & Yang, 2024) also provides an analysis of the RBK algorithm when sampling from a determinantal point process, and other proposals include greedy block Kaczmarz algorithms such as (Liu & Gu, 2021) which require evaluating the complete residual vector at each iteration.
Unfortunately many of these proposals apply only to consistent linear systems, and all of them require at least a preprocessing step in which the entire data matrix must be accessed. Such preprocessing can be prohibitively expensive for very large-scale problems, for which 1) it can be necessary to furnish an approximate solution without processing the entire data set even once (for instance, in the semi-infinite case), and 2) it can be impossible to manipulate more than a tiny subset of the data at a time due to storage constraints. This leads us to the central question of our work:
Can RBK, or some variant thereof, be applied to solve inconsistent linear systems without preprocessing the input matrix?
Viewing the problem (1) as a uniform mixture of $m$ distinct rows, we make the natural choice to focus on the case in which the sampling distribution for the algorithm is also uniform. Our results readily generalize to both weighted and semi-infinite problems of the form
$\min_{x\in\mathbb{R}^{n}}\mathbb{E}_{i\sim\mu}\left[(a_{i}^{\top}x-b_{i})^{2}\right],$ (6)
for which the corresponding algorithms would independently sample $k$ indices $i_{1},\ldots,i_{k}\sim\mu$. This is especially relevant for scientific applications, in which data is often continuous and may be sampled using a physics-based probability distribution. For example, in the case of neural network wavefunctions (Hermann et al., 2023), the sampling distribution is known as the Born probability density.
1.5 Contributions
We demonstrate that the RBK algorithm with uniform sampling (RBK-U) converges in a Monte Carlo sense to a weighted least-squares solution for both consistent and inconsistent linear systems. In particular, Theorem 3.1 shows that convergence is obtained by both expectation values of individual iterates and tail averages of the sequence of iterates. The weight matrix depends on the block size $k$ and the matrix $A$, but not on the vector $b$. Our results provide a new perspective on RBK in the inconsistent case, going beyond previous analyses that only characterized proximity to the ordinary least-squares solution.
Unfortunately, Theorem 3.1 does not guarantee that RBK-U is robust for every problem. To the contrary, when the problem contains many blocks $A_{S}$ that are nearly singular, the condition number of the weight matrix and the variance of the iterates can become arbitrarily large. See Figure 2 for numerical examples that manifest these issues.
One way to overcome these difficulties is to make stronger assumptions on the data. For example, Theorem 3.2 shows that convergence is obtained to the standard least-squares solution when the data arises from certain multivariate Gaussian distributions. Furthermore, in this case both the variance of the iterates and the convergence parameter $\alpha$ can be explicitly bounded. When the singular values of the covariance matrix decay rapidly, the convergence rate can be much faster than mSGD.
Many realistic problems are not well-modeled by Gaussian data. To provide a more general solution, we propose to regularize the RBK iterations as follows:
$x_{t+1}=x_{t}+A_{S_{t}}^{\top}(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}(b_{S_{t}}-A_{S_{t}}x_{t}).$ (7)
The RBK iteration (5) is recovered as $\lambda\rightarrow 0$, but we propose to instead use a small constant $\lambda>0$. This choice corresponds to a stochastic proximal point algorithm with a large, constant step size $\nicefrac{{1}}{{\lambda}}$, which to our knowledge has not been analyzed by previous works. We refer to this algorithm as the regularized block Kaczmarz method, or ReBlocK.
Similar to RBK-U, we show that ReBlocK with uniform sampling (ReBlocK-U) converges in a Monte Carlo sense to a weighted least-squares solution; see Theorem 4.1. Unlike for RBK-U, both the condition number of the weight matrix and the variance of the iterates can be controlled in terms of just $\lambda$ and some coarse properties of the data $A,b$. This makes ReBlocK-U much more reliable than RBK-U in practice. As an added benefit, ReBlocK iterations can be significantly more efficient than RBK iterations; see Sections 4.1 and F.2.1.
Our initial motivation for this work came from the problem of calculating natural gradient directions for deep neural networks. In Section 5, we explain how this setting naturally lends itself to the kinds of linear least-squares solvers we have studied in this paper. Encouragingly, Figure 3 shows that ReBlocK-U outperforms mSGD and RBK-U for this problem. Altogether, our results suggest that ReBlocK can be a more effective algorithm than mSGD for realistic problems that exhibit rapid singular value decay, especially when only moderate accuracy is required.
While our focus is on the case of uniform sampling, the same proof techniques can be applied to other sampling distributions. For example, in Appendix E we show that sampling from an appropriate determinantal point process can in theory enable tail averages of ReBlocK iterates to converge rapidly to the ordinary least-squares solution for inconsistent problems.
1.6 Related Works
Our regularized algorithm, ReBlocK, is closely related to the iterated Tikhonov-Kaczmarz method (De Cezaro et al., 2011). ReBlocK can also be viewed as a specific application of stochastic proximal point algorithms (sPPA) (Bertsekas, 2011; Asi & Duchi, 2019; Davis & Drusvyatskiy, 2019) for solving stochastic optimization problems with objective functions of the least-squares type. To ensure the exact convergence of sPPA, diminishing step sizes are required; see for example (Pătraşcu, 2021). In contrast, our work investigates the convergence of sPPA with a large constant step size for the special case of a least-squares loss. Such an approach allows for aggressive updates throughout the algorithm, potentially improving its practical efficiency.
Our work is also related to the nearly concurrent paper (Dereziński et al., 2025), which introduces Tikhonov regularization into the RBK iterations just like ReBlocK. (Dereziński et al., 2025) focuses on consistent systems that are preprocessed with a randomized Hadamard transform, and the regularization gives rise to optimal convergence rates in the presence of Nesterov acceleration. On the other hand, our work focuses on solving inconsistent systems without any preprocessing step, in which case the regularization is needed to ensure the stability of the algorithm.
The use of Nesterov acceleration represents a much broader trend in the development of stochastic iterative algorithms. Originally proposed by (Nesterov, 1983), this technique has been studied extensively in the context of both stochastic gradient and coordinate descent methods; see for example (Shalev-Shwartz & Zhang, 2013; Allen-Zhu et al., 2016; Jain et al., 2018a; Agarwal et al., 2020). Relative to block projection methods like RBK and ReBlocK, Nesterov acceleration represents an independent and complementary way to improve upon the convergence rate of mSGD. Incorporating Nesterov acceleration into our algorithms is a promising direction for future work.
2 Overview of the Analysis
Our results are stated in terms of a unified framework that includes RBK, ReBlocK, and even mSGD as special cases. Consider the following iteration:
  1. 1.
    Sample $S_{t}\sim\rho$.
  2. 2.
    Update
    $x_{t+1}=x_{t}+A_{S_{t}}^{\top}M(A_{S_{t}})(b_{S_{t}}-A_{S_{t}}x_{t}).$ (8)
Here $M(\cdot):\mathbb{R}^{k\times n}\rightarrow\mathbb{R}^{k\times k}$ takes in the sampled block $A_{S_{t}}$ and returns a positive semidefinite “mass” matrix. RBK is recovered by setting $M(A_{S_{t}})=(A_{S_{t}}A_{S_{t}}^{\top})^{+}$ and ReBlocK by setting $M(A_{S_{t}})=(A_{S_{t}}A_{S_{t}}^{\top}+\lambda|S_{t}|I)^{-1}$. See Algorithm 1 for the full procedure with and without tail averaging.
Once the function $M(A_{s})$ is chosen, let
$W(S)=I_{S}^{\top}M(A_{S})I_{S},\;P(S)=A_{S}^{\top}M(A_{S})A_{S},$ (9)
where $I_{S}$ represents the rows of the $n\times n$ identity matrix whose indices are in $S$. These quantities are natural because they enable the general iteration (8) to be rewritten as
$x_{t+1}=(I-P(S_{t}))x_{t}+A^{\top}W(S_{t})b.$ (10)
Note that $P(S_{t})$ is a projection matrix in the case of RBK.
Next, let
$\overline{W}=\mathbb{E}_{S\sim\rho}\left[W(S)\right],\;\overline{P}=\mathbb{E}_{S\sim\rho}\left[P(S)\right].$ (11)
Additionally, define the weighted solution $x^{(\rho)}$ and the weighted residual $r^{(\rho)}$ via
$x^{(\rho)}=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\left\lVert Ax-b\right\rVert_{\overline{W}}^{2},\;r^{(\rho)}=b-Ax^{(\rho)}.$ (12)
When the solution to the weighted problem is not unique, let $x^{(\rho)}$ refer to the minimal-norm solution.
Using these definitions, we can further rewrite the iteration (8) as
$x_{t+1}-x^{(\rho)}=(I-P(S_{t}))(x_{t}-x^{(\rho)})+A^{\top}W(S_{t})r^{(\rho)}.$ (13)
Since the normal equations for (12) can be written as $A^{\top}\overline{W}r^{(\rho)}=0$, we observe that the final term in (13) vanishes in expectation. Indeed, identifying the appropriate weighted solution $x^{(\rho)}$ to enable the generalized iteration (8) to be written in the form of (13), namely as a linear contraction of the error plus a zero-mean additive term, is the main technical innovation underlying our results. From here, the analysis of (Epperly et al., 2024) can be readily generalized to show that convergence to $x^{(\rho)}$ is obtained.
Algorithm 1 Generalized iterative least-squares solver with optional tail averaging
  Input: Data $A,b$, block size $k$, initial guess $x_{0}$
  Input: Mass matrix $M(A_{S})$, sampling distribution $\rho$
  Input: Total iterations $T$, optional burn-in time $T_{b}$
  for $t=0$ to $T-1$ do
     Sample $S_{t}\sim\rho$
     $x_{t+1}=x_{t}+A_{S_{t}}^{\top}M(A_{S_{t}})(b_{S_{t}}-A_{S_{t}}x)$
  end for
  if $T_{b}$ is not provided then
     Return $x_{T}$
  end if
  $\overline{x}_{T}=\frac{1}{T-T_{b}}\sum_{t=T_{b}+1}^{T}x_{t}$
  Return $\overline{x}_{T}$
3 RBK without Preprocessing
Theorem 3.1.
Consider the RBK-U algorithm, namely Algorithm 1 with $M(A_{S})=(A_{S}A_{S}^{\top})^{+}$ and $\rho=\mathbf{U}(m,k)$. Let $\alpha=\sigma^{+}_{\rm min}(\overline{P})$ and assume that $x_{0}\in\operatorname{range}(A^{\top})$. Then the expectation of the RBK-U iterates $x_{T}$ converges to $x^{(\rho)}$ as
$\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert.$ (14)
Furthermore, the tail averages $\overline{x}_{T}$ converge to $x^{(\rho)}$ as
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq\left(1-\alpha\right)^{T_{b}+1}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}\\ +\frac{1}{\alpha^{2}(T-T_{b})}\;\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}.$ (15)
To our knowledge, this is the first result for inconsistent linear systems that characterizes the exact solution to which the randomized block Kaczmarz iterates converge, albeit in a Monte Carlo sense. The $O(1/T)$ convergence rate for the tail-averaged bound is optimal for row-access methods, and a reasonable default for the burn-in time is $T_{b}=T/2$; see (Epperly et al., 2024) for a more thorough discussion of these points. The proof of Theorem 3.1, which takes advantage of an orthogonal decomposition of the error term $x_{t+1}-x^{(\rho)}$, can be found in Appendix B.
Unfortunately, Theorem 3.1 does not imply robust convergence for general problems. For one, the convergence parameter $\alpha$, which affects every term in the bound, is difficult to analyze in general. Worse, for problems containing nearly singular blocks $A_{S}$, the weight matrix $\overline{W}=\mathbb{E}_{S\sim\rho}\left[I_{S}^{\top}(A_{S}A_{S}^{\top})^{+}I_{S}\right]$ can be arbitrarily poorly conditioned and the variance term $\mathbb{E}\,_{S\sim\rho}||A_{S}^{+}r^{(\rho)}_{S}||^{2}$ can be arbitrarily large. See Figure 2 for numerical examples that manifest these problems.
In the next subsection, we show that these problems can be overcome when the data arises from a Gaussian distribution. A more general solution is provided by the ReBlocK algorithm introduced in Section 4.
3.1 Linear Least-squares with Gaussian Data
Consider a random problem with the data $A,b$ generated as
$\begin{bmatrix}a_{i}^{\top}&b_{i}\end{bmatrix}\sim\mathcal{N}(0,Q)$ (16)
independently for each index $i$, where $Q\in\mathbb{R}^{(n+1)\times(n+1)}$ is a positive semidefinite matrix. Furthermore, let $Q_{n}$ be the top left $n\times n$ block of $Q$, assume for simplicity that $Q_{n}$ is full rank, and let $Q_{n}=LL^{\top}$ be its Cholesky decomposition. Denote the singular values of $L$ by $\sigma_{1}\geq\dots\geq\sigma_{n}>0$. Finally, denote by $y$ the solution to the underlying statistical problem, which is unique since $Q_{n}$ is full rank:
$y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right].$ (17)
Theorem 3.2.
Consider the RBK-U algorithm, namely Algorithm 1 with $M(A_{S})=(A_{S}A_{S}^{\top})^{+}$ and $\rho=\mathbf{U}(m,k)$, applied to the randomly generated finite problem (16). Then the results of Theorem 3.1 hold with
$\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$
Furthermore, as long as $k\geq 6$ and $\operatorname{rank}(L)\geq 2k$, the variance term satisfies
$\lim_{m\rightarrow\infty}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}\leq\frac{200}{\sigma_{2k}^{2}}\cdot\lim_{m\rightarrow\infty}\frac{\left\lVert r\right\rVert^{2}}{m}.$ (18)
Finally, the convergence parameter $\alpha$ satisfies
$\lim_{m\rightarrow\infty}\alpha\geq C_{n,k}\max\left\{\frac{k\sigma_{n}^{2}}{\|L\|_{F}^{2}},\max_{2\leq\ell<k}\frac{(\ell-1)\sigma_{n}^{2}}{\sum_{i\geq k-\ell-1}\sigma_{i}^{2}}\right\}$ (19)
with $C_{n,k}\rightarrow 1$ as $n\rightarrow\infty$ for fixed $k$.
Thus, in the case of Gaussian data and large $m$, the ordinary least-squares solution is recovered, the variance of the iterates is bounded, and the convergence rate $\alpha$ improves at least linearly with the block size $k$. In addition, the following corollary, which is based on Corollary 3.4 of (Dereziński & Rebrova, 2024), shows that RBK converges much faster than mSGD for polynomially decaying singular values.
Corollary 3.3.
Consider the setting of Theorem 3.2 with fixed $k\leq n/2$, and assume the $L$ factor has polynomial spectral decay $\sigma_{i}^{2}\leq i^{-\beta}\sigma_{1}^{2}$ for all $i$ and some $\beta>1$. Then the convergence parameters of RBK and mSGD satisfy
$\lim_{m\rightarrow\infty}\alpha^{\rm RBK}\geq Ck^{\beta}\tfrac{\sigma_{n}^{2}}{\|L\|_{F}^{2}},\;\lim_{m\rightarrow\infty}\alpha^{\rm mSGD}\leq k\tfrac{\sigma_{n}^{2}}{\|L\|_{F}^{2}}$ (20)
for some constant $C=C(\beta)>0$.
Similarly, the faster convergence rate of RBK over mSGD extends to exponentially decaying singular values. The proofs of Theorems 3.2 and 3.3, provided in Appendix B, rely on a connection between RBK-U for Gaussian data and sketch-and-project with Gaussian sketch matrices. These results generalize the techniques of (Derezinski & Mahoney, 2021; Dereziński & Rebrova, 2024) to the case of inconsistent linear systems, improving upon the results of (Rebrova & Needell, 2021) in terms of both the convergence rate and the variance. It is worth noting that Gaussian data is just one way to generate a benign or “incoherent” problem. We expect that Theorems 3.2 and 3.3 could be generalized to other cases such as problems that have been preprocessed using a randomized Hadamard transform, potentially extending the results of (Dereziński & Yang, 2024) to the inconsistent case.
3.2 Implementation Details
To stably implement the RBK iteration (5), we employ an SVD-based least-squares solver to calculate $A_{S_{t}}^{+}(b_{S_{t}}-A_{S_{t}}x).$ The most expensive part of this procedure is the SVD, which has an asymptotic cost of $O(nk^{2})$. This cost is greater than the cost of an mSGD iteration, which is $O(nk)$. Whether it is worthwhile to expend this extra cost depends on the structure of the matrix $A$, and especially on its singular values as discussed in the previous section.
3.3 Numerical Demonstration
We test our theoretical results for the RBK-U and mSGD algorithms on two problems with Gaussian data, with results in Figure 1. We apply tail averaging to each algorithm to observe convergence beyond the variance horizon. As expected, tail-averaged RBK-U (TA-RBK-U) converges much more rapidly than tail-averaged mSGD (TA-mSGD) in the presence of fast singular value decay. Further details on these experiments and a link to the code can found in Appendix F.
Figure 1: Comparison of methods on two problems with Gaussian data, with no singular value decay (Left) and rapid singular value decay (Right). The reported quantity is the suboptimality of the approximate solution $x$ for the unweighted problem (1), namely the value of $\epsilon$ for which $\left\lVert Ax-b\right\rVert=(1+\epsilon)\left\lVert r\right\rVert$. The vertical dotted line indicates the burn-in period of $T_{b}=300$, before which results are shown for individual iterates.
4 Robustness through Regularization
To address the shortcomings of RBK-U in the case of general data, we propose to incorporate regularization into the RBK algorithm. A natural way to do this is to replace the RBK iteration with a stochastic proximal point iteration (Asi & Duchi, 2019), namely
$x_{t+1}=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\left[\left\lVert A_{S_{t}}x-b_{S_{t}}\right\rVert^{2}+\lambda k\left\lVert x-x_{t}\right\rVert^{2}\right].$ (21)
This minimization problem leads to the closed form
$x_{t+1}=x_{t}+A_{S_{t}}^{\top}(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}(b_{S_{t}}-A_{S_{t}}x_{t}),$ (22)
and that the RBK iteration (5) is recovered in the limit $\lambda\rightarrow 0$. To incorporate a mild regularization without significantly slowing down the convergence, we propose to use a small, constant value of $\lambda$ throughout the algorithm. We suggest $\lambda=0.001$ as a practical default value, but to obtain optimal performance the value may need to be tuned on a case-by-case basis; see Section F.1.1 for further discussion. We refer to the resulting scheme as the regularized block Kaczmarz method, or ReBlocK.
Theorem 4.1.
Consider the ReBlocK-U algorithm, namely Algorithm 1 with $M(A_{S})=(A_{S}A_{S}^{\top}+\lambda kI)^{-1}$ and $\rho=\mathbf{U}(m,k)$. Let $\alpha=\sigma_{\rm min}^{+}(\overline{P})$ and assume $x_{0}\in\operatorname{range}(A^{\top})$. Then the expectation of the ReBlocK iterates $x_{T}$ converges to $x^{(\rho)}$ as
$\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert.$ (23)
Furthermore, the tail averages $\overline{x}_{T}$ converge to $x^{(\rho)}$ as
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}\\ +\frac{1}{2\lambda\alpha^{2}(T-T_{b})}\frac{\left\lVert r^{(\rho)}\right\rVert^{2}}{m}.$ (24)
Finally, when the squared row norms of $A$ are uniformly bounded by a constant $N$, then $\kappa(\overline{W})\leq 1+\frac{N}{\lambda}$ and the weighted residual $r^{(\rho)}$ can be bounded in terms of the ordinary least-squares residual $r$:
$\left\lVert r^{(\rho)}\right\rVert^{2}\leq\left(1+\frac{N}{\lambda}\right)\left\lVert r\right\rVert^{2}.$ (25)
Note that the values of $x^{(\rho)}$ and $\alpha$ here differ from those in Section 2 due to the different choice of $M(A_{S})$ used by ReBlocK. To our knowledge, Theorem 4.1 is the first result characterizing the convergence of a stochastic proximal point algorithm with a constant step size for inconsistent linear systems. The advantage relative to RBK-U is that ReBlocK-U is able to control both the variance of the iterates and the conditioning of the weight matrix $\overline{W}$ in terms of the reciprocal of the regularization parameter $\lambda$. As a result, the algorithm converges robustly even when the problem contains many nearly singular blocks $A_{S}$. It is worth noting that these advantages could also be attained by truncating the singular values that are smaller than $\lambda$ in the RBK iteration (5). However, the ReBlocK iteration is additionally justified by the fact that it is cheaper to implement than the RBK iteration; see Section 4.1 for further discussion of this point.
The proof of Theorem 4.1 is provided in Appendix C. Relative to the proof of Theorem 3.1, the proof of Theorem 4.1 is more complicated because the ReBlocK iteration does not lead to an orthogonal decomposition of the error term $x_{t+1}-x^{(\rho)}$. Instead, the proof relies on a bias-variance decomposition inspired by (Défossez & Bach, 2015; Jain et al., 2018b; Epperly et al., 2024), which also leads to the extra factors of $2$ in the convergence bound.
We are not yet able to analyze the convergence rate of ReBlocK-U, even in the case of Gaussian data, as to our knowledge there is no existing work bounding the quality of a regularized Gaussian sketch. However, convergence to the ordinary least-squares solution is obtained for a broader class of noisy linear least-squares problems; see Appendix D. Additionally, a fast rate of convergence to the ordinary least-squares solution can be shown when sampling from an appropriate determinantal point process; see Appendix E.
4.1 Implementation Details
To implement the ReBlocK iterations (7), we directly calculate the $k\times k$ matrix $A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI$ and then use a Cholesky-based linear system solver to calculate $(A_{S_{t}}A_{S_{t}}^{\top}+\lambda kI)^{-1}(b_{S_{t}}-A_{S_{t}}x)$. This computation is stable as long as $\lambda$ is not chosen to be too small. Using this approach, the most expensive part of the ReBlocK iteration is calculating $A_{S_{t}}A_{S_{t}}^{\top}$, which has an asymptotic cost of $O(nk^{2})$ just like RBK. Nonetheless, in practice the matrix-matrix multiplication for ReBlocK can be significantly cheaper than the SVD used for RBK. For example, in the experiments of Section 5, ReBlocK iterations are about five times faster than RBK iterations, as reported in Section F.2.1. For the largest problems, the ReBlocK iterations could be further accelerated using iterative solvers; see for example Section 4.1 of (Dereziński et al., 2025).
4.2 Numerical Demonstration
We now test the ReBlocK-U algorithm on two problems whose columns are discretized representations of continuous functions, leading the matrices $A$ to contain many nearly singular blocks of rows. We set $\lambda=0.001$ and generate the inconsistency using random Gaussian noise. We find that TA-RBK-U is unstable for these problems, as expected. Furthermore, TA-ReBlocK-U converges much faster than TA-mSGD in the presence of rapid singular value decay. Further details on these experiments can be found in Section F.1, and an exploration of the effect of choosing different values of $\lambda$ can be found in Section F.1.1.
Figure 2: Comparison of methods on two inconsistent problems with many nearly singular blocks, with mild singular value decay (Left) and rapid singular value decay (Right). The reported quantity is the suboptimality of the approximate solution $x$ for the unweighted problem (1), namely the value of $\epsilon$ for which $\left\lVert Ax-b\right\rVert=(1+\epsilon)\left\lVert r\right\rVert$. The vertical dotted line indicates the burn-in period of $T_{b}=300$, before which results are shown for individual iterates.
Figure 3: Comparison of methods for calculating natural gradient directions for a small neural network. The network parameters $\theta$ are taken from three snapshots of a single training run, with one snapshot from the “pre-descent” phase before the loss begins to decrease (Left), one snapshot from the “descent” phase during which the loss decreases rapidly (Middle), and one snapshot from the “post-descent” phase when the loss has stopped decreasing significantly (Right). The algorithms are measured in terms of their progress towards reducing the relative residual $\tilde{r}=\left\lVert Jx-[f_{\theta}-f]\right\rVert/\left\lVert f_{\theta}-f\right\rVert$ for the least-squares problem (31), which measures how well the function-space update direction $Jx$ agrees with the function-space loss gradient $f_{\theta}-f$. The burn-in time is set to $T_{b}\approx T/20$ in each case, as indicated by the vertical dotted line.
5 Natural Gradient Optimization
Our original motivation for this work comes from the problem of training deep neural networks using natural gradient descent (Amari, 1998), which is based on an efficient natural gradient induced by a problem-dependent Riemannian metric. Natural gradient descent has been studied extensively in the machine learning community; see (Martens & Grosse, 2015; Ren & Goldfarb, 2019; Martens, 2020). Furthermore, there is increasing evidence that natural gradient methods can improve the accuracy when training neural networks to solve physical equations. See (Müller & Zeinhofer, 2023; Dangel et al., 2024) for applications to physics-informed neural networks and (Pfau et al., 2020; Schätzle et al., 2023) for applications to neural network wavefunctions.
To elucidate the structure of the natural gradient direction, consider the function learning problem
$\min_{\theta}L(\theta),\quad L(\theta)\coloneqq\frac{1}{2}\int_{\Omega}(f_{\theta}(s)-f(s))^{2}{\rm d}s,$ (26)
where $\Omega\subset\mathbb{R}^{d}$ is the domain of the functions, $f:\Omega\rightarrow\mathbb{R}$ is the target function and $f_{\theta}:\Omega\rightarrow\mathbb{R}$ is a function represented by a neural network with parameters $\theta\in\mathbb{R}^{n}$. The standard definition of natural gradient descent for this problem is
$\theta\leftarrow\theta-\eta G_{N},\quad G_{N}\coloneqq F^{-1}\nabla_{\theta}L(\theta),$ (27)
where $\eta$ is the step size, $F$ is the Fisher information matrix
$F=\int_{\Omega}\nabla_{\theta}f_{\theta}(s)\nabla_{\theta}f_{\theta}(s)^{\top}{\rm d}s=J^{\top}J,$ (28)
and the Euclidean gradient $\nabla_{\theta}L(\theta)$ takes the form
$\nabla_{\theta}L(\theta)=\int_{\Omega}\nabla_{\theta}f_{\theta}(s)(f_{\theta}(s)-f(s)){\rm d}s=J^{\top}[f_{\theta}-f].$ (29)
Here $J:\mathbb{R}^{n}\rightarrow\mathbb{R}^{\Omega}$ represents the Jacobian, which is a linear operator from the space of parameters to the space of real-valued functions on $\Omega$. $J^{\top}$ represents the adjoint.
Calculating the natural gradient direction $G_{N}$ using (27) requires a linear solve against the $n\times n$ matrix $F$, which is very challenging in realistic settings when $n\geq 10^{6}.$ This has motivated the development of approximate schemes such as (Martens & Grosse, 2015). An alternative approach is to reformulate $G_{N}$ using the structure of $F$ and $\nabla_{\theta}L(\theta)$:
$G_{N}$ $=(J^{\top}J)^{-1}J^{\top}[f_{\theta}-f]$ (30)
$=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\left\lVert Jx-[f_{\theta}-f]\right\rVert^{2},$ (31)
where the norm in the final expression is the $L_{2}$-norm in function space. This least-squares formulation has been pointed out for example by (Martens, 2020; Chen & Heyl, 2024; Goldshlager et al., 2024), with the work of Chen and Heyl empowering major advances in the field of neural quantum states. A major goal of the current work is to provide a more solid foundation for the development of natural gradient approximations along these lines.
The natural way to access the data when training a neural network is to sample a set of points in the domain $\Omega$ and evaluate the target function $f$, the network outputs $f_{\theta}$, and the network gradients $\nabla_{\theta}f_{\theta}$ at the sampled points. This is precisely equivalent to sampling a small subset of the rows of (31), which motivates the consideration of row-access least-squares solvers for calculating natural gradient directions. Furthermore, both empirical evidence from scientific applications (Park & Kastoryano, 2020; Wang et al., 2022) and theoretical evidence from the literature on neural tangent kernels (Bietti & Mairal, 2019; Ronen et al., 2019; Cao et al., 2019) suggest that $J$ should be expected to exhibit fast singular value decay, motivating the possibility that RBK and ReBlocK could converge rapidly when solving (31). Note that these observations translate straightforwardly from the simple function learning problem (26) to the realistic problems of training physics-informed neural neteworks or neural network wavefunctions.
5.1 Numerical Demonstration
To test our algorithms for calculating natural gradient directions, we train a small neural network to learn a simple function on the unit interval. We take three snapshots from the training process and form the least-squares problem (31) for each. The methods TA-mSGD, TA-RBK-U, and TA-ReBlocK-U are then compared on these problems with results in Figure 3. The computations are performed on an A100 GPU to simulate a deep learning setting, and progress is measured with wall-clock time on the horizontal axis. TA-ReBlocK-U performs best in every case, achieving an accuracy nearly an order of magnitude better than TA-RBK-U in the first snapshot and a full order of magnitude better than TA-mSGD in the second snapshot. Further details on these experiments can be found in Section F.2. For additional context, we examine the singular values of each problem and confirm that they exhibit exponential decay in the top part of the spectrum; see Section F.2.2.
Our results, though preliminary due to their small scale, suggest that ReBlocK is a promising method for calculating natural gradient directions. This provides justification for the Kaczmarz-inspired SPRING algorithm (Goldshlager et al., 2024) and suggests a family of related methods to be explored by future works. Open questions include how many ReBlock iterations to run between each update of $\theta$, how to incorporate averaging, and how to tune $\lambda$.
6 Conclusions
In this work, we have explored the problem of solving large-scale linear least-squares problems without preprocessing the data matrix $A$. Our results suggest that ReBlocK may be a more effective algorithm than mSGD for inconsistent problems that exhibit rapid singular value decay. More broadly, our work highlights the value of incorporating regularization as a path towards broadening the applicability of block Kaczmarz methods, and our analysis suggests that a Monte Carlo perspective can be useful in elucidating the behavior of randomized block row-access methods in general. Finally, our work provides motivation and suggests new directions for least-squares based natural gradient optimizers.
Acknowledgements
G.G. acknowledges support from the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112. This effort was supported by the SciAI Center, and funded by the Office of Naval Research (ONR), under Grant Number N00014-23-1-2729 (J.H., L.L.). L.L. is a Simons Investigator in Mathematics. This research used the Savio computational cluster resource provided by the Berkeley Research Computing program at the University of California, Berkeley. We thank Ethan Epperly, Robert Webber, and Ruizhe Zhang for their thoughtful discussions and feedback.
Disclaimer
This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.
Appendix A Helpful Lemmas
In this appendix we prove five lemmas which will help us analyze both RBK and ReBlocK. Lemma A.1 shows how the expectation of the iterates evolves. Lemma A.2 shows that the iterates of the algorithms always stay within $\operatorname{range}(A^{\top})$, and Lemma A.3 provides a contraction property of applying $I-\overline{P}$ to the certain types of vectors. Lemma A.4 shows how to derive bounds on the expected square error of tail-averaged iterates under relevant assumptions. Finally, Lemma A.5 bounds the attained residual based on the condition number of $\overline{W}$.
Lemma A.1 (Convergence of single-iterate expectation).
Consider the generalized iteration (8) with some fixed choice of a positive semidefinite mass matrix $M(A_{S})$ and sampling distribution $\rho$, and fix two indices $r<s$. Then the expectation of $x_{s}$ conditioned on $x_{r}$ satisfies
$\mathbb{E}\left[x_{s}-x^{(\rho)}|x_{r}\right]=(I-\overline{P})^{s-r}(x_{r}-x^{(\rho)}).$ (32)
Proof of Lemma A.1.
Fix $t\in r,\ldots,s-1$, let $P_{t}=P(S_{t})$ and $W_{t}=W(S_{t})$. Recall from (13) that the generalized iteration (8) can be reformulated as
$x_{t+1}-x^{(\rho)}=(I-P_{t})(x_{t}-x^{(\rho)})+A^{\top}W_{t}r^{(\rho)}.$ (33)
Recall also that the normal equations for $x^{(\rho)}$ imply that $A^{\top}\overline{W}r^{(\rho)}=0$. Taking the expectation over the choice of $S_{t}$, it thus obtains
$\mathbb{E}\left[x_{t+1}-x^{(\rho)}|x_{t}\right]=(I-\overline{P})(x_{t}-x^{(\rho)})+A^{\top}\overline{W}r^{(\rho)}=(I-\overline{P})(x_{t}-x^{(\rho)}).$ (34)
Using the law of total expectation and the linearity of expectation, iterating this result for $t=r,\ldots,s-1$ yields the lemma. ∎
Lemma A.2.
Consider the generalized iteration (8) with some fixed choice of a positive semidefinite mass matrix $M(A_{S})$ and sampling distribution $\rho$. Then $x^{*},x^{(\rho)}\in\operatorname{range}(A^{\top})$ and if $x_{0}\in\operatorname{range}(A^{\top})$, then $x_{t}\in\operatorname{range}(A^{\top})$ for all $t\geq 0$.
Proof.
Recall that $x^{*}$ is the minimal-norm solution to (1) and $x^{(\rho)}$ is the minimal-norm solution to (12). Suppose for the sake of contradiction that $x^{*}\notin\operatorname{range}(A^{\top})$. Then let $\tilde{x}$ be the projection of $x^{*}$ onto $\operatorname{range}(A^{\top})$. It follows that $\left\lVert\tilde{x}\right\rVert<\left\lVert x^{*}\right\rVert$ and $Ax^{*}=A\tilde{x}$, making $\tilde{x}$ a minimizer of (1) with a smaller norm than $x^{*}$ (contradiction). The same argument holds for $x^{(\rho)}$.
To show $x_{t}\in\operatorname{range}(A^{\top})$ for all $t\geq 0$ we apply an inductive argument. The claim holds for $t=0$ by assumption, so now suppose that it holds for some arbitrary $t\geq 0$. Recall
$x_{t+1}=x_{t}+A_{S_{t}}^{\top}M(A_{S_{t}})(b_{S_{t}}-A_{S_{t}}x_{t}).$ (35)
Then $x_{t}\in\operatorname{range}(A^{\top})$ by assumption and $A_{S_{t}}^{\top}M(A_{S_{t}})(b_{S_{t}}-A_{S_{t}}x_{t})\in\operatorname{range}(A^{\top})$ since $\operatorname{range}(A_{S_{t}}^{\top})\subseteq\operatorname{range}(A^{\top})$. It follows that $x_{t+1}\in\operatorname{range}(A^{\top})$, completing the proof.
Lemma A.3.
Suppose that $x\in\operatorname{range}(\overline{P})$, $\overline{P}\preceq I$, and $\alpha=\sigma_{\rm min}^{+}(\overline{P})$. Then for any $s\geq 0$,
$x^{\top}(I-\overline{P})^{s}x\leq(1-\alpha)^{s}\left\lVert x\right\rVert^{2}$ (36)
and
$\left\lVert(I-\overline{P})^{s}x\right\rVert\leq(1-\alpha)^{s}\left\lVert x\right\rVert.$ (37)
Proof.
First expand $x$ in the basis of eigenvectors of the symmetric matrix $\overline{P}$, then note that every eigenvector that has a nonzero coefficient in the expansion has its eigenvalue in the interval $[\alpha,1]$.
Lemma A.4 (Convergence of tail-averaged schemes).
Consider the generalized iteration (8) with some fixed mass matrix $M(A_{S})$ and sampling distribution $\rho$. Let $\alpha=\sigma_{\rm min}^{+}(\overline{P})$ and suppose that $x_{0}\in\operatorname{range}(A^{\top})$, $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$, and $\overline{P}\preceq I$. Additionally, suppose the stochastic iterates $x_{0},x_{1},\ldots$ satisfy
$\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{t}B+V$ (38)
for all $t$ and some constants $B,V$. Then the tail averages $\overline{x}_{T}$ of the iterates, with burn-in time $T_{b}$, satisfy
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{1}{\alpha(T-T_{b})}V.$ (39)
Proof.
We follow closely the Proof of Theorems 1.2 and 1.3 of (Epperly et al., 2024). Decompose the expected mean square error as
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}=\frac{1}{(T-T_{b})^{2}}\sum_{r,s=T_{b}+1}^{T}\mathbb{E}\left[(x_{r}-x^{(\rho)})^{\top}(x_{s}-x^{(\rho)})\right].$ (40)
For $T_{b}\leq r<s$ bound the covariance term using Lemma A.1 and our assumptions on $\overline{P}$:
$\mathbb{E}\left[(x_{r}-x^{(\rho)})^{\top}(x_{s}-x^{(\rho)})\right]$ $=\mathbb{E}\left[(x_{r}-x^{(\rho)})^{\top}\mathbb{E}\left[x_{s}-x^{(\rho)}|x_{r}\right]\right]$
$=\mathbb{E}\left[(x_{r}-x^{(\rho)})^{\top}(I-\overline{P})^{s-r}(x_{r}-x^{(\rho)})\right]$
$\leq(1-\alpha)^{s-r}\mathbb{E}\,\left\lVert x_{r}-x^{(\rho)}\right\rVert^{2}$
$\leq(1-\alpha)^{s}B+(1-\alpha)^{s-r}V,$
where the third line uses Lemmas A.2 and A.3 and the assumptions $\operatorname{range}{\overline{P}}=\operatorname{range}(A^{\top})$ and $\overline{P}\preceq I$.
Using the coarse bound $(1-\alpha)^{s}B\leq(1-\alpha)^{T_{b}+1}B$ for $s\geq T_{b}+1$, it follows
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{V}{(T-T_{b})^{2}}\sum_{r,s=T_{b}+1}^{T}(1-\alpha)^{|s-r|}.$ (41)
Apply another course bound
$\sum_{r,s=T_{b}+1}^{T}(1-\alpha)^{|s-r|}\leq 2\sum_{r=T_{b}+1}^{T}\sum_{s=0}^{\infty}(1-\alpha)^{s}=\frac{T-T_{b}}{\alpha}$ (42)
to obtain the final result:
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{T_{b}+1}B+\frac{V}{\alpha(T-T_{b})}.$ (43)
Lemma A.5 (Residual bound).
Assuming $\overline{W}\succ 0$, it holds
$\left\lVert r^{(\rho)}\right\rVert^{2}\leq\kappa(\overline{W})\left\lVert r\right\rVert^{2}.$ (44)
Proof.
Use operator norms and the optimality of $x^{(\rho)}$ for the weighted least-squares problem:
$\left\lVert Ax^{(\rho)}-b\right\rVert^{2}$ $\leq\left\lVert\overline{W}^{-1}\right\rVert\left\lVert Ax^{(\rho)}-b\right\rVert_{\overline{W}}^{2}$
$\leq\left\lVert\overline{W}^{-1}\right\rVert\left\lVert Ax^{*}-b\right\rVert_{\overline{W}}^{2}$
$\leq\left\lVert W^{-1}\right\rVert\left\lVert\overline{W}\right\rVert\left\lVert r\right\rVert^{2}$
$=\kappa(\overline{W})\left\lVert r\right\rVert^{2}.$
Appendix B Proofs of RBK Convergence Theorems
In this section we provide the proofs of Theorems 3.1, 3.2 and 3.3. We first prove the following lemma, which can be viewed as a more general version of Theorem 1.2 of (Needell & Tropp, 2014):
Lemma B.1.
Consider the RBK iterates (5) and fix some sampling distribution $\rho$. Suppose that $x_{0}\in\operatorname{range}(A^{\top})$ and $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. Then for all $t$ it holds
$\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq(1-\alpha)^{t}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{1}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}.$ (45)
Proof.
Let $P_{s}=P(S_{s})$ and $W_{s}=W(S_{s})$. Using (13) and the definition $M(A_{S})=(A_{S}A_{S}^{\top})^{+}$, the RBK iteration (5) can be reformulated as
$x_{s+1}-x^{(\rho)}=(I-P_{s})(x_{s}-x^{(\rho)})+A_{S_{s}}^{+}r^{(\rho)}.$ (46)
This represents an orthogonal decomposition of $x_{s+1}-x^{(\rho)}$ since $I-P_{s}$ is the projector onto the null space of $A_{S_{s}}$, which is in turn orthogonal to the range of the pseudoinverse $A_{S_{s}}^{+}$.
It thus holds
$\left\lVert x_{s+1}-x^{(\rho)}\right\rVert^{2}$ $=\left\lVert(I-P_{s})(x_{s}-x^{(\rho)})\right\rVert^{2}+\left\lVert A_{S_{s}}^{+}r^{(\rho)}_{S_{s}}\right\rVert^{2}$
$=(x_{s}-x^{(\rho)})^{\top}(I-P_{s})(x_{s}-x^{(\rho)})+\left\lVert A_{S_{s}}^{+}r^{(\rho)}_{S_{s}}\right\rVert^{2},$
where the second line uses the idempotency of the projector $I-P_{s}$. Note that $x_{s}-x^{(\rho)}\in\operatorname{range}(A^{\top})$ by Lemma A.2 and $\overline{P}\preceq I$ since $P(S)$ is always a projection matrix. Taking expectations and applying Lemma A.3 thus yields
$\mathbb{E}\,\left\lVert x_{s+1}-x^{(\rho)}\right\rVert^{2}$ $=(x_{s}-x^{(\rho)})^{\top}(I-\overline{P})(x_{s}-x^{(\rho)})+\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}$
$\leq(1-\alpha)\mathbb{E}\,\left\lVert x_{s}-x^{(\rho)}\right\rVert^{2}+\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}.$
Iterating from $s=0$ to $s=t-1$ and utilizing $\sum_{s=0}^{t-1}(1-\alpha)^{s}<\sum_{s=0}^{\infty}(1-\alpha)^{s}=\nicefrac{{1}}{{\alpha}}$ yields the desired result. ∎
Proof of Theorem 3.1.
The result follows from the appropriate application of Lemmas A.1, A.4 and B.1. To apply these lemmas it is first required to verify $\overline{P}\preceq I$ and $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. For the first result, it suffices to observe that $P(S)$ is the orthogonal projector onto the row space of $A_{S}$ and thus $P(S)\preceq I$ for all $S$.
For the second result, first note that since $\overline{P}=A^{\top}\overline{W}A$ it is clear that $\operatorname{range}(\overline{P})\subseteq\operatorname{range}(A^{\top})$. Utilizing this containment and the fact that $\overline{P}$ is symmetric, it suffices to show that there is no $x\in\operatorname{range}(A^{\top})$ such that $\overline{P}x=0$. Now, consider any such $x\in\operatorname{range}(A^{\top})$. There must exist a row index $i$ such that $a_{i}^{\top}x\neq 0$. Thus
$x^{\top}\overline{P}x=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[x^{\top}A_{S}^{\top}(A_{S}A_{S}^{\top})^{-1}A_{S}x\right]\geq\frac{k}{m}\cdot\frac{(a_{i}^{\top}x)^{2}}{\left\lVert a_{i}\right\rVert^{2}}>0,$ (47)
where the intermediate step uses the facts that row $i$ is chosen with probability $k/m$ and that $A_{S}^{\top}(A_{S}A_{S}^{\top})^{-1}A_{S}\succeq a_{i}(a_{i}^{\top}a_{i})^{-1}a_{i}^{\top}$ when $i\in S$. It follows that $\overline{P}x\neq 0$ and so indeed $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$.
With the assumptions verified, Lemma A.1 can be applied using $r=0$, $s=T$ to yield
$\mathbb{E}\left[x_{T}\right]-x^{(\rho)}=(I-\overline{P})^{T}(x_{0}-x^{(\rho)}).$ (48)
This directly implies
$\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert$ (49)
using Lemmas A.2 and A.3 and the proven properties of $\overline{P}$.
Furthermore, Lemma B.1 holds and provides the conditions for Lemma A.4 with $B=\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}$, $V=\frac{1}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{+}r^{(\rho)}_{S}\right\rVert^{2}$. The application of Lemma A.4 then directly implies the convergence result for the tail-averaged RBK-U algorithm. ∎
Proof of Theorem 3.2.
In the case of Gaussian data, suppose we are given $m$ data points $\{a_{i},b_{i}\}_{i=1}^{m}$. Then the dataset $A,b$ can be viewed as a finite sample of the underlying statistical linear regression problem
$y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right].$ (50)
Furthermore, the solution $x^{*}$ is the corresponding (minimal-norm) empirical risk minimizer. It is a standard result in asymptotic statistics that as the number of samples $m$ grows, the empirical minimizer $x^{*}$ converges to the true solution $y$ (Lehmann & Casella, 2006; Van der Vaart, 2000).
Now, the RBK-U algorithm is applied by uniformly selecting $k$ samples per iteration from the finite data $A,b$. Nonetheless, in the limiting case where $m\to\infty$, the sampling distribution of $k$ rows from the finite dataset $\{a_{i},b_{i}\}_{i=1}^{m}$ converges to the distribution obtained by sampling $k$ items independently from $\mathcal{N}(0,Q)$. Thus, to obtain results about the RBK-U algorithm in the limit $m\rightarrow\infty$, we can focus on the case in which the $k$ data points $\{a_{1},b_{1}\},\dots,\{a_{k},b_{k}\}$ are directly sampled from $\mathcal{N}(0,Q).$
Denote by $Q=VV^{\top}$ the Cholesky decomposition of $Q$. Let $V=\begin{pmatrix}L\\ d^{\top}\end{pmatrix}$, where :::: and $d^{\top}$ are the top $n$ rows and the last row of $V$, respectively. Recalling that $Q_{n}$ represents the top left $n\times n$ block of $Q$, it follows that $Q_{n}=LL^{\top}$ is the Cholesky decomposition of $Q_{n}$, as previously defined.
In each iteration of RBK-U, the $k$ data points $[a_{1}^{\top}\;b_{1}],\dots,[a_{k}^{\top}\;b_{k}]$ from $\mathcal{N}(0,Q)$ can be equivalently redistributed as $z_{1}^{\top}[L^{\top}\;d],\dots,z_{k}^{\top}[L^{\top}\;d]$ for $z_{i}\sim\mathcal{N}(0,I_{n+1})$. Collecting the random vectors $z_{1},\ldots,z_{k}$ into the columns of a single matrix $Z_{t}$, the RBK-U update can be written as
$x_{t+1}=x_{t}+(Z_{t}^{\top}L^{\top})^{+}(Z_{t}^{\top}d-Z_{t}^{\top}L^{\top}x_{t}).$ (51)
Now, note that $\mathbb{E}\left[a_{i}a_{i}^{\top}\right]=LL^{\top},\mathbb{E}\left[b_{i}a_{i}\right]=Ld$ and $\mathbb{E}\left[b_{i}^{2}\right]=d^{\top}d$. Plugging these identities into the definition of $y$, we find
$y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{[a_{i}^{\top}\;b_{i}]\sim\mathcal{N}(0,Q)}\left[(a_{i}^{\top}x-b_{i})^{2}\right]=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\|L^{\top}x-d\|^{2}.$ (52)
We can thus define an “underlying” residual vector by $\tilde{r}=L^{\top}y-d$. Additionally defining $P_{t}=(Z_{t}^{\top}L^{\top})^{+}Z_{t}^{\top}L^{\top}$, the update (51) can be reformulated as
$x_{t+1}-y=(I-P_{t})(x_{t}-y)+(Z_{t}^{\top}L^{\top})^{+}Z_{t}^{\top}\tilde{r}.$ (53)
Noting from the definition that $\tilde{r}$ is orthogonal to the rows of $L$, we now show $\mathbb{E}\left[(Z_{t}^{\top}L^{\top})^{+}(Z_{t}^{\top}\tilde{r})\right]=0$. In fact, decompose $Z_{t}=Z_{t}^{1}+Z_{t}^{2}$ where $Z_{t}^{1}\in\operatorname{range}(L^{\top})$ and $Z_{t}^{2}\perp\operatorname{range}(L^{\top})$. It follows that $Z_{t}^{1}$ and $Z_{t}^{2}$ are independent, zero-mean Gaussian vectors and
$\mathbb{E}\left[(Z_{t}^{\top}L^{\top})^{+}(Z_{t}^{\top}\tilde{r})\right]=\mathbb{E}\left[((Z_{t}^{1})^{\top}L^{\top})^{+}(Z_{t}^{2})^{\top}\tilde{r}\right]=\mathbb{E}\left[((Z_{t}^{1})^{\top}L^{\top})^{+}\right]\mathbb{E}\left[(Z_{t}^{2})^{\top}\tilde{r}\right]=0.$
We note that a similar independence lemma has also been established in Lemma 3.14 of (Rebrova & Needell, 2021). Furthermore, taking the expectation on both sides of (53), we have
$\mathbb{E}\left[x_{t+1}-x^{*}\right]=(I-\overline{P})(x_{t}-x^{*}),$
where $\overline{P}:=\mathbb{E}\left[P_{t}\right]$. Consequently, in the infinite limit $\mathbb{E}\left[x_{t}\right]$ converges to $y$ using the same logic as the proof of Theorem 3.1, which proves that
$\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$ (54)
In addition, the variance term $\left\lVert A_{S_{t}}^{\top}r_{S_{t}}\right\rVert^{2}$ in Theorem 3.1 takes the form
$\left\lVert(Z^{\top}L^{\top})^{+}Z^{\top}\tilde{r}\right\rVert^{2}.$ (55)
We can thus bound
$\left\lVert(Z^{\top}L^{\top})^{+}Z^{\top}\tilde{r}\right\rVert^{2}=\left\lVert(Z_{1}^{\top}L^{\top})^{+}Z_{2}^{\top}\tilde{r}\right\rVert^{2}\leq\sigma_{\rm min}(Z_{1}^{\top}L^{\top})^{-2}\left\lVert Z_{2}^{\top}\tilde{r}\right\rVert^{2}.$ (56)
We can bound the expectation of the smallest singular value of $Z_{1}^{\top}L^{\top}$ using a similar technique to the proof of Lemma 22 in (Dereziński & Rebrova, 2024). To start, let $L^{\top}=W\Sigma Y^{\top}$ be the SVD of $L^{\top}$. Then
$\sigma_{\rm min}(Z_{1}^{\top}L^{\top})^{2}=\sigma_{\rm min}(Z_{1}^{\top}L^{\top}LZ_{1})=\sigma_{\rm min}(Z_{1}^{\top}W\Sigma^{2}W^{\top}Z_{1}).$ (57)
Letting $W_{2k}$ denote the first $2k$ columns of $W$, and noting that $\Sigma\succeq{\rm diag}(\sigma_{2k},\dots,\sigma_{2k},0,\dots,0)$, it holds
$\sigma_{\rm min}(Z_{1}^{\top}L^{\top})^{2}\geq\sigma_{2k}^{2}\cdot\sigma_{\rm\min}(Z_{1}^{\top}W_{2k}W_{2k}^{\top}Z_{1})=\sigma_{\min}(Z_{1}^{\top}W_{2k}).$
Since the columns of $W_{2k}$ are orthonormal the random matrix $Z_{1}^{\top}W_{2k}$ can be redistributed as a single Gaussian random matrix $G_{2k}$ of size $k\times(2k)$ and with standard normal entries. By Lemma 3.16 of (Rebrova & Needell, 2021), we thus have for $k\geq 6$,
$\mathbb{E}\left[\sigma_{\rm min}^{-2}(Z_{1}^{\top}L^{\top})\right]\leq\frac{20}{(\sqrt{2k}-\sqrt{k})^{2}\sigma_{2k}^{2}}\leq\frac{200}{k\sigma_{2k}^{2}}.$ (58)
Hence, by the independence between $(Z^{\top}L^{\top})^{+}$ and $Z^{\top}\tilde{r}$, we have
$\mathbb{E}\,\|(Z^{\top}L^{\top})^{+}Z^{\top}\tilde{r}\|^{2}\leq\mathbb{E}\left[\sigma_{\min}^{-2}(Z_{1}^{\top}L^{\top})\right]\mathbb{E}\,\|Z_{2}^{\top}\tilde{r}\|^{2}\leq\frac{200}{\sigma_{2k}^{2}}\|\tilde{r}\|^{2}.$
Noting that $\lim_{m\rightarrow\infty}\frac{\|r\|^{2}}{m}=\mathbb{E}\left[(a_{i}^{\top}x^{*}-b_{i})^{2}\right]=\|L^{\top}x^{*}-d\|^{2}=\|\tilde{r}\|^{2}$, we prove (18).
It remains only to bound the value of $\alpha$. Note that the entries of $Z_{t}\in\mathbb{R}^{d\times m}$ are independently drawn from the standard Gaussian distribution.
This property enables us to leverage existing results from (Dereziński & Rebrova, 2024) concerning the spectrum of the matrix $\overline{P}$. Specifically, $P_{t}=(Z_{t}L^{\top})^{+}Z_{t}L^{\top}$, which has a similar formula to the matrix $P$ considered in Equation (6) of (Dereziński & Rebrova, 2024). Then, applying Theorem 3.1 of (Dereziński & Rebrova, 2024) gives the result stated in Theorem 3.2. ∎
Proof of Corollary 3.3..
By Corollary 3.4 of (Dereziński & Rebrova, 2024), if $L$ has the polynomial spectral decay of order $\beta>1$, i.e., $\sigma_{i}^{2}\leq i^{-\beta}\sigma_{1}^{2}$ for all $i$, the dependence of $C_{n,k}$ on $n$ and $k$ in Theorem 3.2 can be eliminated when $k\leq n/2$. Furthermore, there is a constant $C=C(\beta)$ such that for any $k\leq n/2$, the linear convergence rate satisfies
$\lim_{m\rightarrow\infty}\alpha^{\rm RBK}\geq C\frac{k^{\beta}\sigma_{n}^{2}}{\|L\|_{F}^{2}}.$
Regarding the convergence rate of mSGD, let us consider solving a fixed number of samples, say $m$, of $\{a_{i}^{\top},b_{i}\}$. It is demonstrated in (Needell et al., 2014) that, with a minibatch size of 1 and importance sampling, the corresponding convergence parameter can be at most $\alpha^{\rm SGD}\leq 1/\kappa_{\rm dem}(\frac{1}{m}A_{m}A_{m}^{\top})$ with $A_{m}=[a_{1},\dots,a_{m}]$ for convergent step sizes. When a larger minibatch size $k$ is used, (Jain et al., 2018b) shows that the learning rate can be increased at most linearly, corresponding to a rate of at most $\alpha^{\rm mSGD}\leq k/\kappa_{\rm dem}(\frac{1}{m}A_{m}A_{m}^{\top})$. It follows by the random matrix theory (Johnstone, 2001; Tao, 2012) that $\lim_{m\rightarrow\infty}\frac{1}{m}A_{m}A_{m}^{\top}=LL^{\top}$. Hence, when $m$ goes to infinity, the linear convergence rate of mSGD satisfies
$\lim_{m\rightarrow\infty}\alpha^{\rm mSGD}\leq\frac{k\sigma_{n}^{2}}{\|L\|_{F}^{2}}.$
Appendix C Proofs of ReBlocK Convergence Theorems
This appendix contains the proof of the ReBlocK convergence bound Theorem 4.1. We first establish a useful lemma which bounds the error of the individual ReBlocK iterates under arbitrary sampling distributions.
Lemma C.1.
Consider the ReBlocK iterates (7) and fix some sampling distribution $\rho$. Suppose that $x_{0}\in\operatorname{range}(A^{\top})$ and $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. Then for all $t$ it holds
$\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{t}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{1}{2\lambda k\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$ (59)
Proof.
Since the ReBlocK iteration does not admit a simple orthogonal decomposition, we instead proceed by utilizing a bias-variance decomposition inspired by (Défossez & Bach, 2015; Jain et al., 2018b; Epperly et al., 2024).
Let $P_{s}=P(S_{s})$ and $W_{s}=W(S_{s})$. Bias and variance sequences are defined respectively by
$d_{0}=x_{0}-x^{(\rho)},\;d_{s+1}=(I-P_{s})d_{s},$ (60)
$v_{0}=0,\;v_{s+1}=(I-P_{s})v_{s}+A^{\top}W_{s}r^{(\rho)}.$ (61)
Intuitively, the bias sequence captures the error due to the initialization $x_{0}\neq x^{(\rho)}$ and the variance sequence captures the rest of the error. It can be verified by mathematical induction and (13) that $x_{s}-x^{(\rho)}=d_{s}+v_{s}$ for all $s$. As a result, it holds
$\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq 2\mathbb{E}\,\left\lVert d_{t}\right\rVert^{2}+2\mathbb{E}\,\left\lVert v_{t}\right\rVert^{2}.$ (62)
Note also that it is a simple extension of Lemma A.2 that $d_{s},v_{s}\in\operatorname{range}(A^{\top})$ for all $s$.
To analyze the bias, first calculate
$\mathbb{E}\left[\left\lVert d_{s+1}^{2}\right\rVert|d_{s}\right]=d_{s}^{\top}\mathbb{E}\left[(I-P_{s})^{2}\right]d_{s}=d_{s}^{\top}(I-2\overline{P}+\mathbb{E}_{S\sim\rho}\left[P(S)^{2}\right])d_{s}.$ (63)
Observe that
$P(S)=A_{S}^{\top}(A_{S}A_{S}^{\top}+k|S|I)^{-1}A_{S}\preceq A_{S}^{\top}(A_{S}A_{S}^{\top})^{-1}A_{S}\preceq I,$ (64)
and hence $P(S)^{2}\preceq P(S)$ and $\mathbb{E}_{S\sim\rho}\left[P(S)^{2}\right]\preceq\overline{P}$. As a result, $I-2\overline{P}+\mathbb{E}_{S\sim\rho}\left[P(S)^{2}\right]\preceq I-\overline{P}$. Additionally leveraging the properties $d_{s}\in\operatorname{range}(A^{\top})=\operatorname{range}(\overline{P})$ and $\overline{P}\preceq I$, Lemma A.3 implies
$\mathbb{E}\left[\left\lVert d_{s+1}^{2}\right\rVert|d_{s}\right]\leq(1-\alpha)\mathbb{E}\,\left\lVert d_{s}\right\rVert^{2}.$ (65)
Iterating this inequality yields a simple bound on the bias term:
$\mathbb{E}\,\left\lVert d_{t}\right\rVert^{2}\leq(1-\alpha)^{t}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}.$ (66)
The analysis of the variance term is less straightforward. To begin, note that $v_{s}$ follows the same recurrence as $x_{s}-x^{(\rho)}$, so we can apply a slight modification of Lemma A.1 to the variance sequence to obtain
$\mathbb{E}\left[v_{s}\right]=(I-\overline{P})^{s}v_{0}=0.$ (67)
Now, square the iteration for $v_{s+1}$ conditioned on $v_{s}$:
$\mathbb{E}\left[\left\lVert v_{s+1}\right\rVert^{2}|v_{s}\right]$ $=v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))^{2}\right]v_{t}+2v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))A^{\top}W(S)r^{(\rho)}\right]+\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2}$
$\leq(1-\alpha)\left\lVert v_{s}\right\rVert^{2}+2v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))A^{\top}W(S)r^{(\rho)}\right]+\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2},$
where the second step uses our previous observations that $\mathbb{E}_{S\sim\rho}\left[(I-P(S))^{2}\right]\preceq I-\overline{P}$. $v_{s}\in\operatorname{range}(\overline{P})$, and Lemma A.3. Take expectations over $v_{s}$ as well to eliminate the cross-term, yielding
$\mathbb{E}\,\left\lVert v_{s+1}\right\rVert^{2}\leq(1-\alpha)\mathbb{E}\,\left\lVert v_{s}\right\rVert^{2}+\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2}.$
Iterating this last inequality for $s=0,\ldots,t-1$, using $\sum_{s=0}^{t-1}(1-\alpha)^{s}<\sum_{s=0}^{\infty}(1-\alpha)^{s}=\nicefrac{{1}}{{\alpha}}$, and recalling that $v_{0}=0$, we find
$\mathbb{E}\,\left\lVert v_{t}\right\rVert^{2}\leq\frac{1}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2}.$ (68)
Now, note that
$A^{\top}W(S)r^{(\rho)}=A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}r^{(\rho)}_{S}.$ (69)
Applying the singular value decomposition of $A_{S}$ and some elementary calculus, it is verified that regardless of the singular values of $A_{S}$,
$\left\lVert A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}\right\rVert\leq\frac{1}{2\sqrt{\lambda k}}.$ (70)
The variance bound can thus be simplified as
$\mathbb{E}\,\left\lVert v_{t}\right\rVert^{2}\leq\frac{1}{4\lambda k\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$ (71)
Combining the bias bound (66) and the variance bound (71) using (62) yields the desired result. ∎
Proof of Theorem 4.1.
The result will follow from the appropriate application of Lemmas A.1, A.4 and C.1. To apply these lemmas it is first required to verify y $\overline{P}\preceq I$ and $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. For the first result, it suffices to observe that $P(S)\preceq P_{RBK}(S)\preceq I$ for all $S$.
For the second result, the logic follows almost identically to the same part of the proof of Theorem 3.1. The only difference is in how we show that $\overline{P}x\neq 0$ for any $x\in\operatorname{range}(A^{\top})$. Like before, we start by noting that there must exist a row index $i$ such that $a_{i}^{\top}x\neq 0$. We then bound
$x^{\top}\overline{P}x=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[x^{\top}A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}A_{S}x\right]\geq\frac{k}{m}\cdot\frac{(a_{i}^{\top}x)^{2}}{\left\lVert AA^{\top}+k\lambda I\right\rVert_{2}}>0,$ (72)
where the intermediate step uses the facts that $(A_{S}A_{S}^{\top}+k\lambda I)^{-1}\succeq(AA^{\top}+k\lambda I)^{-1}$, that row $i$ is chosen with probability $k/m$, and that $x^{\top}A_{S}^{\top}A_{S}x\geq x^{\top}a_{i}a_{i}^{\top}x$ when $i\in S$. It follows that $\overline{P}x\neq 0$ and so indeed $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$.
With the assumptions verified, Lemma A.1 can be applied using $r=0$, $s=T$ to yield
$\mathbb{E}\left[x_{T}\right]-x^{(\rho)}=(I-\overline{P})^{T}(x_{0}-x^{(\rho)}).$ (73)
This directly implies
$\left\lVert\mathbb{E}\left[x_{T}\right]-x^{(\rho)}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{(\rho)}\right\rVert.$ (74)
using Lemmas A.2 and A.3 and the proven properties of $\overline{P}.$
The conditions of Lemma C.1 are also satisfied, the results of which provide the conditions for Lemma A.4 with $B=\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}$, $V=\frac{1}{2\lambda k\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$ Lemma A.4 then leads to the following bound for the tail averages:
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x^{(\rho)}\right\rVert^{2}+\frac{1}{2\lambda k\alpha^{2}(T-T_{b})}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$ (75)
Since $\rho=\mathbf{U}(m,k)$, the last term can be simplified using $\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}=k\left\lVert r^{(\rho)}\right\rVert^{2}/m$. The desired tail-averaged convergence bound follows.
The next step is to bound the condition number $\kappa(\overline{W})$, given that $\left\lVert a_{i}\right\rVert^{2}\leq N$ for all $i$. First note that the row norm bound implies
$\left\lVert A_{S}A_{S}^{\top}\right\rVert=\left\lVert A_{S}^{\top}A_{S}\right\rVert=\left\lVert\sum_{i\in S}a_{i}a_{i}^{\top}\right\rVert\leq\sum_{i\in S}\left\lVert a_{i}a_{i}^{\top}\right\rVert\leq Nk.$ (76)
As a result,
$\lambda kI\preceq A_{S}A_{S}^{\top}+\lambda kI\preceq(N+\lambda)kI.$ (77)
Now, when $S$ is sampled uniformly, $\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[I_{S}^{\top}I_{S}\right]=\frac{k}{m}I$. It follows that
$\overline{W}=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[I_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}I_{S}\right]\succeq\frac{1}{(N+\lambda)m}I,$ (78)
$\overline{W}=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[I_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}I_{S}\right]\preceq\frac{1}{\lambda m}I.$ (79)
Putting these together implies
$\kappa(\overline{W})\leq 1+\frac{N}{\lambda},$ (80)
and the final claim of the theorem follows by Lemma A.5. ∎
Appendix D Noisy Linear Least-squares
In this section we consider the special case in which the inconsistency in the problem is generated by zero-mean noise. This case is more general than the case of Gaussian data, but less general than the case of arbitrary inconsistency. We define such a problem to have each row independently generated as
$a_{i}\sim\mathbf{A},\;z_{i}\sim\mathbf{Z},\;b_{i}=a_{i}^{\top}y+z_{i},$ (81)
where $\mathbf{A}$ is arbitrary, $\mathbf{Z}$ satisfies $\mathbb{E}_{z\sim\mathbf{Z}}\left[z\right]=0$, and $y\in\mathbb{R}^{n}$. This scenario might arise, for example, if the entries of $b$ are calculated via noisy measurements of $y$ along the directions $a_{i}$.
For such problems with increasing numbers of rows, both RBK and ReBlocK converge (in a Monte Carlo sense) to the ordinary least-squares solution rather than a weighted solution. This can be viewed as an advancement over previous results on noisy linear systems, such as (Needell, 2010), which show only convergence to a finite horizon even for arbitrarily large noisy systems. The advance comes from treating the noise model explicitly and applying tail averaging to the algorithm.
Theorem D.1.
Consider the RBK-U algorithm, namely Algorithm 1 with $M(A_{S})=(A_{S}A_{S})^{+}$ and $\rho=\mathbf{U}(m,k)$, applied to the noisy linear least-squares problem (81). Then the results of Theorem 3.1 hold and
$\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$ (82)
Theorem D.2.
Consider the ReBlocK-U algorithm, namely Algorithm 1 with $M(A_{S})=(A_{S}A_{S}^{\top}+\lambda kI)^{-1}$ and $\rho=\mathbf{U}(m,k)$, applied to the noisy linear least-squares problem (81). Then the results of Theorem 4.1 hold and
$\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$ (83)
Unified proof of Theorems D.1 and D.2.
Our strategy will be to first show that $x^{*}$ converges to $y$, then show that $x^{(\rho)}$ converges to $x^{*}$.
For the first part, note that for fixed $m$, the dataset $A,b$ can be viewed as a finite sample of the underlying statistical linear regression problem
$y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{a\sim\mathbf{A},z\sim\mathbf{Z}}\left[(a^{\top}(x-y)-z)^{2}\right].$ (84)
Furthermore, the ordinary least-squares solution $x^{*}$ is the corresponding empirical risk minimizer. It is a standard result in asymptotic statistics that as the number of samples $m$ grows, the empirical minimizer $x^{*}$ converges to the true solution $y$ (Lehmann & Casella, 2006; Van der Vaart, 2000).
For the second part, recall that the weighted solution $x^{(\rho)}$ is characterized by
$\overline{P}x^{(\rho)}=A^{\top}\overline{W}b=\overline{P}x^{*}+A^{\top}\overline{W}r.$ (85)
It has been verified in the proofs of Theorems 3.1 and 4.1 that $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. Combining this with the results of Lemma A.2 implies that $x^{*},x^{(\rho)}\in\operatorname{range}(\overline{P})$. Applying the pseudoinverse of $\overline{P}^{+}$ on the left thus yields
$x^{(\rho)}=x^{*}+\overline{P}^{+}A^{\top}\overline{W}r,$ (86)
and it then suffices to show that
$\lim_{m\rightarrow\infty}\overline{P}^{+}A^{\top}\overline{W}r=0.$ (87)
As $m\rightarrow\infty$, the distribution of $k$ rows sampled uniformly at random from $A$ converges to the distribution of $k$ rows sampled independently from $\mathbf{A}$. Similarly, the distribution of $k$ entries of the residual $r=Ax^{*}-b$ sampled uniformly from $A,b$ converges to the distribution of $k$ values sampled independently from $\mathbf{Z}$. By the noisy construction, these residual values are independent from each other as well as from the corresponding rows of $A$. Call the resulting distributions $\mathbf{A}^{k}$ and $\mathbf{Z}^{k}$. It then holds
$\lim_{m\rightarrow\infty}\overline{P}=\lim_{m\rightarrow\infty}\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[P(A_{S})\right]=\mathbb{E}_{A_{S}\sim\mathbf{A}^{k}}\left[P(A_{S})\right]\coloneqq\hat{P},$ (88)
$\lim_{m\rightarrow\infty}\overline{P}^{+}A^{\top}\overline{W}r=\hat{P}^{+}\lim_{m\rightarrow\infty}\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[A_{S}^{\top}M(A_{S})r_{S}\right]=\hat{P}^{+}\mathbb{E}_{A_{S}\sim\mathbf{A}^{k}}\left[A_{S}^{\top}M(A_{S})\right]\mathbb{E}_{r_{S}\sim\mathbf{Z}^{k}}\left[r_{S}\right]=0.$ (89)
It follows that
$\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$ (90)
Appendix E Sampling from a Determinantal Point Process
In Section 4, we have presented the convergence of ReBlocK to a weighted least-squares solution under the uniform sampling distribution $\rho=\mathbf{U}(m,k)$. However, the explicit convergence rate $1-\alpha$ is still unknown. Motivated by recent work on Kaczmarz algorithms with determinantal point process (DPP) sampling (Dereziński & Yang, 2024), we now show that under DPP sampling, ReBlocK converges to the ordinary least-squares solution and the convergence of ReBlocK can have a much better dependence on the singular values of $A$ than mSGD.
The sampling distribution that we consider is $\rho=k\text{-}{\rm DPP}(AA^{\top}+\lambda kI)$, namely
$\Pr[S]=\frac{\det(A_{S}A_{S}^{\top}+\lambda kI)}{\sum_{|S^{\prime}|=k}\det(A_{S^{\prime}}A_{S^{\prime}}^{\top}+\lambda kI)}.$ (91)
Our analysis is similar to parts of the nearly concurrent work (Dereziński et al., 2025), though (Dereziński et al., 2025) does not incorporate a fixed size $k$ for their regularized DPP distribution..
Theorem E.1.
Consider the ReBlocK algorithm with k-DPP sampling, namely $M(A_{S})=(A_{S}A_{S}^{\top}+\lambda kI)^{-1}$ and $\rho=k\text{-}{\rm DPP}(AA^{\top}+\lambda kI)$. Let $\alpha=\sigma_{\rm min}^{+}(\overline{P})$ and assume $x_{0}\in\operatorname{range}(A^{\top})$. Then the expectation of the ReBlocK iterates $x_{T}$ converges to $x^{*}$ as
$\left\lVert\mathbb{E}\left[x_{T}\right]-x^{*}\right\rVert\leq(1-\alpha)^{T}\left\lVert x_{0}-x^{*}\right\rVert.$ (92)
Furthermore, the tail averages $\overline{x}_{T}$ converge to $x^{*}$ as
$\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{*}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x_{0}-x^{*}\right\rVert^{2}\\ +\frac{1}{2\lambda\alpha^{2}(T-T_{b})}\cdot\max_{i}(r_{i}^{2}).$ (93)
Moreover, for any $\ell<k$ the convergence parameter $\alpha$ satisfies
$\alpha\geq\frac{k-\ell}{(k-\ell)+\kappa_{\ell}^{2}(A)+(m+k-2\ell)\frac{\lambda}{\sigma_{n_{r}}^{2}}},$ (94)
where $\kappa_{\ell}^{2}(A):=\sum_{j>\ell}^{n_{r}}\sigma_{j}^{2}(A)/\sigma_{n_{r}}^{2}(A)$ and $\sigma_{1}(A)\geq\dots\geq\sigma_{n_{r}}(A)>0=\sigma_{n_{r+1}}(A)=\dots=\sigma_{n}(A)$ are the singular values of $A$.
By choosing $\ell=1$ and a small value of $\lambda$, the above theorem indicates that when $n_{r}=n$ the value of $\alpha$ is at least on the order of $k/\kappa^{2}_{\rm dem}(A)$ where $\kappa_{\rm dem}:=\|A\|_{F}/\sigma_{\min}(A)$ is the Demmel condition number. To understand the meaning of the bound more generally, suppose additionally that $\left\lVert a_{i}\right\rVert=1$ for all $i$ so that $\left\lVert A\right\rVert_{F}^{2}=m$. Then, choosing $\ell=k/2$ we can rewrite (94) as
$\begin{split}\alpha&\geq\frac{k}{k+2\kappa_{k/2}^{2}(A)+2m\lambda/\sigma_{n}^{2}}\\ &\approx\frac{k}{2\kappa^{2}_{k/2}(A)+2\lambda\kappa_{\rm dem}^{2}(A)}.\end{split}$ (95)
As noted in Section 3.1, the convergence parameter for mSGD is bounded by
$\alpha^{\rm SGD}\leq k/\kappa_{\rm dem}^{2}(A).$
Hence, when the singular values of $A$ decay rapidly and $\lambda$ is significantly smaller than one, ReBlocK with DPP sampling can have a much faster convergence rate than mSGD.
Our bound for the variance of the tail-averaged estimator is crude and could likely be improved using techniques such as those of (Dereziński et al., 2025). Nonetheless, it is interesting to note that the dependence of the variance on the residual vector can be worse in the case of DPP sampling than in the case of uniform samping. This is because the DPP distribution can potentially sample the residual vector in unfavorable ways. It is thus unclear whether DPP sampling is a good strategy for arbitrary inconsistent systems, even if it can be implemented efficiently
Proof.
The majority of the proof follows similarly to the proofs of Theorems 3.1 and 4.1, and we only highlight the differences, of which there are three: the fact that $x^{(\rho)}=x^{*}$, the value of the parameter $\alpha$, and the variance term in the tail-averaged bound.
We begin by verifying that $x^{(\rho)}=x^{*}$. For a vector $u\in\mathbb{R}^{m}$, we denote its elementary symmetric polynomial by $p_{k}(u):=\sum_{S\in\binom{[m]}{k}}\prod_{i\in S}u_{i}$, where $\binom{[m]}{k}$ is the set of all subsets of size $k$ from the set $[m]=\{1,\ldots,m\}$. Applying equation (5.3) of (Dereziński & Yang, 2024) by using $B=AA^{\top}+\lambda I$ leads to
$\overline{W}=\mathbb{E}_{S\sim\rho}\left[I_{S}^{\top}(I_{S}(AA^{\top}+\lambda I)I_{S}^{\top})^{-1}I_{S}\right]=\frac{U{\rm diag}(p_{k-1}(q_{-1}),\ldots,p_{k-1}(q_{-m}))U^{\top}}{p_{k}(q)},$ (96)
where $A=U\Sigma V^{\top}$ denotes the compact singular value decomposition of $A$, and $q_{i}=\sigma_{i}^{2}+\lambda$ when $i\leq n_{r}$ and $q_{i}=\lambda$ otherwise, with $\sigma_{1}\geq\dots\geq\sigma_{n_{r}}$ representing the sorted singular values and $n_{r}$ the rank of $A$. Denote by $\overline{W}=UDU^{\top}$ with the diagonal matrix $D:=\frac{{\rm diag}(p_{k-1}(q_{-1}),\ldots,p_{k-1}(q_{-m}))}{p_{k}(q)}\succ 0$. Then, by the definition $x^{(\rho)}={\rm argmin}_{x\in\mathbb{R}^{n}}\|Ax-b\|_{\overline{W}}^{2}$, we have
$A^{\top}\overline{W}Ax^{(\rho)}=A^{\top}\overline{W}b.$
Note that $A^{\top}\overline{W}A=V\Sigma D\Sigma V^{\top}$ and $A^{\top}\overline{W}=V\Sigma DU^{\top}b$. Thus, we can apply $V^{\top}$ on the left and use $D\succ 0$ to conclude
$\Sigma^{2}V^{\top}x^{(\rho)}=\Sigma U^{\top}b.$
Applying $V$ now on the left implies $A^{\top}Ax^{(\rho)}=A^{\top}b$, i.e., $x^{(\rho)}\in{\rm argmin}_{x\in\mathbb{R}^{n}}\|Ax-b\|^{2}$. Since $x^{(\rho)}\in\operatorname{range}(A^{T})$ this implies that $x^{(\rho)}=x^{*}$.
Regarding $\alpha$, plugging (96) into the definition of $\overline{P}$ leads to
$\overline{P}=A^{\top}\overline{W}A=V\Sigma^{\top}\frac{{\rm diag}(p_{k-1}(q_{-1}),\ldots,p_{k-1}(q_{-m}))}{p_{k}(q)}\Sigma V^{\top}.$ (97)
We now follow closely the logic of the proof of Lemma 4.1 of (Dereziński & Yang, 2024), making appropriate modifications to handle the fact that in GRK we invert $AA^{\top}+\lambda I$ rather than just $AA^{\top}$. Note also that through (97) we have explicitly verified that $\operatorname{range}(\overline{P})=\operatorname{range}(A^{T})$, which is the needed condition for the contraction properties to hold in our convergence bound.
First, for any $\ell<k$ we can construct an approximation matrix
$B_{\ell}=U{\rm diag}\left(q_{1},\ldots q_{m}\right)U^{\top}+\frac{1}{k-\ell}\sum_{j>\ell}^{m}q_{j}I.$ (98)
Note that we have used a mildly simpler and looser approximation than (Dereziński & Yang, 2024) by replacing $\frac{k-\ell-1}{k-\ell}$ with $1$ in the first term.
The same arguments as in (Dereziński & Yang, 2024) then imply that
$\alpha=\sigma_{\min}^{+}(\overline{P})\geq\sigma_{\min}^{+}(A^{\top}B_{\ell}^{-1}A).$ (99)
Furthermore we have
$\sigma_{\min}^{+}(A^{\top}B_{\ell}^{-1}A)$ $=\sigma_{\min}^{+}\left(V\Sigma^{\top}U^{\top}B_{\ell}^{-1}U\Sigma V^{\top}\right)$
$=\sigma_{\min}\left(V{\rm diag}\left(\frac{\sigma_{1}^{2}}{q_{1}+\frac{1}{k-\ell}\sum_{j>\ell}^{m}q_{j}},\ldots,\frac{\sigma_{n_{r}}^{2}}{q_{n_{r}}+\frac{1}{k-\ell}\sum_{j>\ell}^{m}q_{j}}\right)V^{\top}\right)$
$=\frac{\sigma_{n_{r}}^{2}}{q_{n_{r}}+\frac{1}{k-\ell}\sum_{j>\ell}^{m}q_{j}}$
$=\frac{\sigma_{n_{r}}^{2}}{\sigma_{n_{r}}^{2}+\frac{1}{k-\ell}\sum_{j>\ell}^{n_{r}}\sigma_{j}^{2}+\frac{m+k-2\ell}{k-\ell}\lambda}$
$=\frac{k-\ell}{(k-\ell)+\kappa_{\ell}^{2}(A)+(m+k-2\ell)\frac{\lambda}{\sigma_{n_{r}}^{2}}}.$
This completes the bound for $\alpha$.
Finally, we consider the variance term in the tail-averaged bound. Following the proofs of Lemmas C.1 and 4.1, we can obtain a variance term of
$\frac{1}{2\lambda k\alpha^{2}(T-T_{b})}\mathbb{E}\,_{S\sim\rho}\left\lVert r_{S}\right\rVert^{2}$ (100)
which is esentially the same as for ReBlocK-U but with the ordinary least-squares residual $r$ instead of the weighted residual $r^{(\rho)}$. Unfortunately, in the case of DPP sampling, we have no guarantee on how $S$ samples the residual vector $r^{(\rho)}$. Thus, the best we can do is uniformly bound $\left\lVert r_{S}\right\rVert^{2}\leq k\cdot\max_{i}(r_{i}^{2})$. This yields the bound in the theorem. ∎
Appendix F Details of Numerical Experiments
In this appendix we provide more details on the numerical examples throughout the paper. The code to run the experiments can be found at https://github.com/ggoldsh/block-kaczmarz-without-preprocessing.
F.1 Experiments with random matrices
The experiments in Sections 3.3 and 4.2 differ only in how the matrix $A$ is generated. For the left panel in Figure 1, we generate the matrix $A$ by setting each entry independently as $A_{ij}\sim\mathcal{N}(0,1)$. For the right panel, we construct $A=GU$ where $G\in\mathbb{R}^{m\times n}$ has independent standard normal entries and and $U$ has $\sigma_{i}=1/i^{2}$ and random orthonormal singular vectors. The condition number of the matrix $A$ is $\kappa(A)\approx 1$ in the left panel and $\kappa(A)\approx 10^{4}$ in the right panel.
The examples in Figure 2 are constructed as discretizations of underlying continuous problems, which is both a realistic scenario and serves to ensure that $A$ contains many nearly singular blocks. In both cases, we first generate an $n\times n$ matrix $C$. For the left panel we set $C=I$, whereas for the right panel we set $C$ with singular values $\sigma_{i}=1/i^{2}$ and random orthonormal singular vectors. Once $C$ is constructed, we define a set of $n$ functions $f_{1},\ldots,f_{n}$ by
$f_{j}(s)=\sum_{\ell=1}^{n}C_{j\ell}T_{\ell}(s),$ (101)
where $T_{\ell}$ is the $\ell^{\rm th}$ Chebyshev polynomial of the first kind. The functions $f$ correspond to the columns of $A$.
Next, we construct a vector $v$ of $m$ coordinates uniformly spaced across the interval $[-1,1]$, namely
$v_{i}=\left(-1+2\frac{i-1}{m-1}\right).$ (102)
Finally, the matrix $A$ is designed as
$A_{ij}=f_{j}(v_{i}),$ (103)
so that column $j$ of $A$ is a dicretized representation of the function $f_{j}$. The condition number of the resulting matrices $A$ are $\kappa(A)\approx 11$ in the left panel and $\kappa(A)\approx 3.6\cdot 10^{4}$ in the right panel.
For all experiments in Sections 3.3 and 4.2 we set $m=10^{5}$, $n=10^{2}$, and $k=30$. Furthermore, the vectors $b$ are constructed by setting
$b_{i}=a_{i}^{\top}y+z_{i}$ (104)
where $y\sim\mathcal{N}(0,I_{n})$ and $z_{i}\sim\mathcal{N}(0,10^{-4})$. The initial guess is $x_{0}=0$ and the number of iterations is $T=10^{4}$, which is equivalent to three passes over the data. The step size for minibatch SGD is constant within each run and has been tuned independently for each example, specifically to be as large as possible without introducing any signs of instability, up to a factor of $2$. The value of $\lambda$ for ReBlocK has been set to $\lambda=0.001$ and is not optimized on a per-example basis. The reported quantity is the suboptimality of the solution for the unweighted problem (1), namely $\epsilon=\left\lVert Ax-b\right\rVert/\left\lVert Ax^{*}-b\right\rVert-1=\left\lVert Ax-b\right\rVert/\left\lVert r\right\rVert-1$. Since it is expensive to calculate the residual, this quantity is only calculated and reported every $10$ iterations. These experiments with random matrices are carried out in double precision on a 1.1 GHz Quad-Core Intel Core i5 CPU.
F.1.1 Effect of Regularization Parameter $\lambda$
In Figure 4, we investigate the effect of the parameter $\lambda$ on the performance of ReBlocK. The problem is the same as in the right panel of Figure 2, in which $A$ contains many ill-conditioned blocks $A_{S}$ and exhibits rapid singular value decay. It is observed that choosing $\lambda=1$ approximately reproduces the results of mSGD, while $\lambda=1e-3,1e-6,1e-9$ converges much faster, with $\lambda=1e-9$ beginning to show signs of instability. The results suggest that the performance of ReBlocK is only moderately sensitive to the choice of $\lambda$, since the parameter must be varied by multiple orders of magnitude to drastically change the performance.
Figure 4: Performance of TA-ReBlocK on the problem from the right panel of Figure 2 with various values of $\lambda$. The reported quantity is the suboptimality of the approximate solution $x$ for the unweighted problem (1), namely the value of $\epsilon$ for which $\left\lVert Ax-b\right\rVert=(1+\epsilon)\left\lVert r\right\rVert$. The vertical dotted line indicates the burn-in period of $T_{b}=300$, before which results are shown for individual iterates.
F.2 Natural Gradient Experiments
In this section we describe in detail the numerical experiments of Section 5. The training problem for the neural network is a simple function regression task on the unit interval. The target function is chosen to be periodic to avoid any consideration of boundary effects, and the neural network is correspondingly designed to be periodic by construction. Specifically, the target function is constructed as
$f(s)=q(\sin(2\pi s)),$ (105)
with the polynomial $q$ defined as
$q(s)=\frac{1}{\sqrt{d}}\sum_{\ell=1}^{d}c_{\ell}T_{\ell}(s)$ (106)
for $d=15$ and each $c_{\ell}$ chosen randomly from a standard normal distribution. The resulting function is pictured in Figure 5.
Figure 5: Target function for neural network training.
The neural network model is a simple periodic ResNet (He et al., 2016) with 5 layers. For input $s\in\mathbb{R}$, the network outputs $f(s)\in\mathbb{R}$ given by
$y_{1}=\tanh(W_{1}\sin(2\pi s)+b_{1}),$ (107)
$y_{i}=y_{i-1}+\tanh(W_{i}y_{i-1}+b_{i});\;i=2,3,4,$ (108)
$f(s)=W_{5}y_{4}+b_{5}.$ (109)
The intermediate layers have dimensions $y_{i}\in\mathbb{R}^{50}$ and the weight matrices $W_{i}$ and bias vectors $b_{i}$ have the appropriate dimensions to match. The weights are initialized using a Lecun normal initialization and the biases are all initialized to zero. The parameters are collected into a single vector $\theta\in\mathbb{R}^{7801}$ for convenience, leading to the neural network function $f_{\theta}(r)$.
The loss function is defined as in (26) and the network is trained using subsampled natural gradient descent, as described for example in (Ren & Goldfarb, 2019), with a batch size of $N_{b}=200$, a Tikhonov regularization of $\lambda=0.01$, and a step size of $\eta=0.5$. This corresponds to the parameter update
$\theta_{t+1}=\theta_{t}-\eta J_{S}^{\top}(J_{S}J_{S}^{\top}+N_{b}\lambda I)^{-1}[f_{\theta}-f]_{S},$ (110)
where $S$ represents the set of $N_{b}$ sample points. The setting of $N_{b}=200$ is intended to represent something close to a full batch training regime, which is only practical since we are studying a very small network on a very compact, low-dimensional domain. The resulting training curve is presented in Figure 6.
Figure 6: Training curve for the neural network function regression example. The three vertical lines indicate the training snapshots that are used to generate the least-squares problems studied in Figure 3.
From this training run, three snapshots are taken and used to generate the least-squares problems for Figure 3, following Equation 31. The first snapshot is from the “pre-descent” phase before the loss begins to decrease, the second is from the “descent” phase during which the loss decreases rapidly, and the third is from the “post-descent” phase when the loss has stopped decreasing significantly. The snapshots are indicated by the vertical dotted lines in Figure 6. The algorithms TA-mSGD, TA-RBK, and TA-ReBlocK are then tested on the resulting problems wth an initial guess of $x_{0}=0$ and a batch size of $k=30$. The batch size of $k=30$ is meant to represent the realistic scenario when each iteration uses too few samples to thoroughly represent the target function. Furthermore, the continuous problems are treated directly by uniformly sampling $k=30$ points from the domain $[0,1]$ at each iteration and calculating the network outputs and gradients at these points.
The step size for minibatch SGD is constant within each run and has been tuned independently for each example, specifically to be as large as possible without introducing any signs of instability, up to a factor of $2$. The value of $\lambda$ for ReBlocK has been set to $\lambda=0.001$ and is not optimized on a per-example basis. The reported quantity is the relative residual $\tilde{r}=\left\lVert Jx-[f_{\theta}-f]\right\rVert/\left\lVert f_{\theta}-f\right\rVert$, which measures how well the function-space update direction $Jx$ agrees with the function-space loss gradient $f_{\theta}-f$. This quantity is estimated accurately by sampling $m=2\cdot 10^{4}$ points from $[0,1]$. Furthermore, since it is expensive to calculate this residual, the quantity is only calculated and reported at approximately $10^{3}$ evenly spaced iterations for each run of each algorithm.
F.2.1 Iteration Cost of Each Algorithm
In Figure 7, we report the number of iterations per second for mSGD, RBK, and ReBlocK for the experiments of Section 5. We find that RBK iterations are approximately five times slower than ReBlocK iterations and six times slower than mSGD iterations for these particular problems. Interestingly, there is only a slight difference in speed between the ReBlocK and mSGD iterations. The experiments are conducted in single precision on an A100 GPU to simulate a deep learning setting.
Figure 7: Iteration cost of each algorithm for the natural gradient experiments of Section 5.
F.2.2 Singular Value Analysis
In Figure 8, we analyze the singular values of the Jacobian operator $J$ for each experiment in Figure 3. We first approximate $J$ by a finite matrix $A$ by sampling $m=2\cdot 10^{4}$ points from the domain $[0,1]$, then use SciPy’s svds function with the ARPACK solver to approximate the top $200$ singular values of $A$. We find that in every case the top singular values decay exponentially up to some threshold, after which the tail decays more slowly. The initial exponential decay helps to explain the faster convergence rate of ReBlocK relative to mSGD.
Figure 8: Top $200$ singular values of the input matrix $A$ for each natural gradient-based least-squares problem. The left, middle, and right panels correspond to the same panels in Figure 3, and the vertical dotted line represents the batch size of $k=30$ used in the least-squares solvers for Figure 3.